OpenStudy (anonymous):

A Particle moves along a straight line so that it's velocity, v ms^-1 at time t seconds is given by v=6e^(3t) +4. When t = 0, the displacement, s, of the particle is 7 metres. Find an expression for s in terms of t. I know im supposed to use integration but im a bit confused on how, can someone please explain how to do it. Thanks!

OpenStudy (anonymous):

\[\int 6e^{3t}+4 dt\]

OpenStudy (anonymous):

2 e^3t+ 4t +5

OpenStudy (cathyangs):

go imranmeah91! :D I agree with what little knowledge I know. did you type this up? XD or is particle supposed to be capitalized?

OpenStudy (ash2326):

we are given v=6e^(3t)+4 v=dx/dt ,x=displacement dx/dt=6e^(3t)+4 for getting a relation of x , we have to integrate the above eq. on integration we get x=6e^(3t)/3+4t+c c=constant of integration x=2e^3t+4t+c it's given when t=0 x=7 put t=0 7=2+c so c=5 x=2e^3t+4t+5

OpenStudy (anonymous):

Is the equation for integrating it not x^n+1 /n +1, so where is the plus one part?

OpenStudy (ash2326):

but here we don't have any t term, and integral of constant like 4 is 4t or u can say 4t^0---->4t^(0+1)/(0+1)=4t

OpenStudy (anonymous):

How come 4 becomes 4t?

OpenStudy (anonymous):

4 = 4 t^0

OpenStudy (ash2326):

lau 4 can also be written as 4t^(0) integral of 4t^0=4t^(0+1)/(0+1) i.e 4t